# A Comprehensive Survey on Benchmarks and Solutions in Software Engineering of LLM-Empowered Agentic System

### v1

@@ Abstract
The integration of Large Language Models (LLMs) into software engineering has catalyzed a paradigm shift from traditional rule-based systems to sophisticated agentic systems capable of autonomous problem-solving. Despite this transformation, the field lacks a comprehensive understanding of how benchmarks and solutions interconnect, hindering systematic progress and evaluation. This survey presents the first holistic analysis of LLM-empowered software engineering, bridging the critical gap between evaluation methodologies and solution approaches. We systematically analyze 150+ recent papers and organize them into a comprehensive taxonomy spanning two major dimensions: (1) Solutions, categorized into prompt-based, fine-tuning-based, and agent-based paradigms, and (2) Benchmarks, covering code generation, translation, repair, and other specialized tasks. Our analysis reveals how the field has evolved from simple prompt engineering to complex agentic systems incorporating planning and decomposition, reasoning and self-refinement, memory mechanisms, and tool augmentation. We present a unified pipeline that illustrates the complete workflow from task specification to final deliverables, demonstrating how different solution paradigms address varying complexity levels across software engineering tasks. Unlike existing surveys that focus on isolated aspects, we provide full-spectrum coverage connecting 50+ benchmarks with their corresponding solution strategies, enabling researchers to identify optimal approaches for specific evaluation criteria. Furthermore, we identify critical research gaps and propose actionable future directions, including multi-agent collaboration frameworks, self-evolving code generation systems, and integration of formal verification with LLM-based methods. This survey serves as a foundational resource for researchers and practitioners seeking to understand, evaluate, and advance LLM-empowered software engineering systems.
@@ I Introduction
Selection Papers of This Survey: To ensure comprehensive coverage and high quality, we systematically collected recent papers from multiple sources including top-tier conferences (NeurIPS, ICML, ICLR, ACL, EMNLP, ICSE, FSE, ASE), journals (TSE, TOSEM, TPAMI), and a small amount of preprint servers (arXiv, OpenReview). Our selection criteria include: (1) Papers published between 2023 and 2025 that explicitly address LLM-based approaches for software engineering; (2) Works that contribute novel benchmarks, datasets, or evaluation metrics; (3) Studies proposing new methods, architectures, or techniques for LLM-empowered agentic systems; (4) Papers with significant empirical evaluation demonstrating practical impact. We initially collected over 500 papers and applied these criteria to select 150+ high-quality papers that form the foundation of this survey. We prioritize papers that have been peer-reviewed at reputable venues while also including influential preprints that have garnered significant attention from the research community.
@@ VII Challenges and Future Directions
VII-A Scalability: From Token Limits to Hierarchical Cognition
@@ VII-B Evaluation: From Functional Correctness to Production Readiness
VII-C Domain Adaptation: From General Patterns to Specialized Expertise
@@
VII-D Multi-Agent Coordination: From Individual Capability to Collective Intelligence
@@
VII-E Continuous Learning: From Static Models to Evolving Expertise
@@
VII-F Responsible Deployment: From Ethical Quagmires to Trustworthy Collaboration

### v2

@@ Abstract
The integration of Large Language Models (LLMs) into software engineering has driven a transition from traditional rule-based systems to autonomous agentic systems capable of solving complex problems. However, systematic progress is hindered by a lack of comprehensive understanding of how benchmarks and solutions interconnect. This survey addresses this gap by providing the first holistic analysis of LLM-powered software engineering, offering insights into evaluation methodologies and solution paradigms. We review over 150 recent papers and propose a taxonomy along two key dimensions: (1) Solutions, categorized into prompt-based, fine-tuning-based, and agent-based paradigms, and (2) Benchmarks, including tasks such as code generation, translation, and repair. Our analysis highlights the evolution from simple prompt engineering to sophisticated agentic systems incorporating capabilities like planning, reasoning, memory mechanisms, and tool augmentation. To contextualize this progress, we present a unified pipeline illustrating the workflow from task specification to deliverables, detailing how different solution paradigms address various complexity levels. Unlike prior surveys that focus narrowly on specific aspects, this work connects 50+ benchmarks to their corresponding solution strategies, enabling researchers to identify optimal approaches for diverse evaluation criteria. We also identify critical research gaps and propose future directions, including multi-agent collaboration, self-evolving systems, and formal verification integration. This survey serves as a foundational guide for advancing LLM-driven software engineering. We maintain a GitHub repository that continuously updates the reviewed and related papers at https://github.com/lisaGuojl/LLM-Agent-SE-Survey.
@@ I Introduction
Selection Papers of This Survey: To ensure comprehensive coverage and high quality, we systematically collected recent papers from multiple sources including top-tier conferences (NeurIPS, ICML, ICLR, ACL, EMNLP, ICSE, FSE, ASE), journals (TSE, TOSEM, TPAMI), and a small amount of preprint servers (arXiv, OpenReview). Our selection criteria include: (1) Papers published between 2023 and 2025 that explicitly address LLM-based approaches for software engineering; (2) Works that contribute novel benchmarks, datasets, or evaluation metrics; (3) Studies proposing new methods, architectures, or techniques for LLM-empowered agentic systems; (4) Papers with significant empirical evaluation demonstrating practical impact.
@@ VII Challenges and Future Directions
While the integration of LLMs has catalyzed significant advances in software engineering, the path toward fully autonomous and reliable agentic systems is fraught with fundamental challenges. These obstacles, spanning scalability, evaluation, and responsible deployment, are not merely incremental hurdles but rather define the critical research frontiers for the field. This section outlines these principal challenges and, in turn, proposes corresponding future research directions poised to overcome them, shaping the next generation of software engineering AI.
VII-A Scalability: From Token Limits to Hierarchical Cognition
A primary challenge is one of scale, where the architectural constraints of current LLMs prevent meaningful engagement with real-world software complexity. While context windows have expanded, they remain insufficient for enterprise systems where codebases routinely exceed millions of lines of code. This limitation manifests as project amnesia, where agents lose track of high-level architectural patterns and design conventions during extended tasks. Current mitigation strategies, such as vector database retrieval, provide only partial solutions, as they fetch isolated snippets of context without comprehending the intricate relationships that define a software architecture. At this scale, the computational economics of iterative refinement also become prohibitive, hindering the application of agents to complex, system-wide tasks.
To overcome these scalability limitations, future research must pivot from expanding linear context windows to developing hierarchical cognitive architectures for code understanding. This necessitates a paradigm shift from processing code as a flat sequence of tokens to reasoning over structured, multi-modal representations like Code Property Graphs (CPGs). Such architectures must be coupled with advanced tiered memory systems, such as combining a volatile short-term memory for immediate context with a persistent long-term memory storing architectural knowledge and project conventions. By learning to navigate these compressed, structured representations, agents can transcend token limits to reason about system-wide properties, marking a crucial transition from agents that merely write code to agents that truly comprehend software.
The field currently faces an evaluation crisis, where prevailing benchmarks and metrics incentivize superficial correctness over the qualities that define production-ready code. An over-reliance on pass@k on synthetic, function-level datasets like HumanEval creates a significant reality gap, failing to measure critical non-functional requirements such as security, performance, and maintainability. Consequently, agents are optimized to hack the benchmark rather than to produce robust, high-quality software. Recent analyses reveal that a majority of algorithmically successful solutions on benchmarks like SWE-Bench would fail in production due to the introduction of new security flaws, performance regressions, or violations of coding standards. Without evaluation frameworks that mirror real software lifecycle concerns, progress remains illusory.
Addressing this evaluation gap requires a move toward holistic, lifecycle-aware benchmarking. The next generation of evaluation frameworks should be built around dynamic, digital twin environments that simulate the entire software development lifecycle. Within these sandboxed environments, agents must be assessed on a suite of production-oriented criteria, including the introduction of security vulnerabilities, performance regressions, and the accumulation of technical debt. Furthermore, evaluation must incorporate economic and operational metrics, such as the computational cost of a solution and the human effort required for review. Developing such comprehensive benchmarks will close the gap between academic research and industrial practice, ensuring that progress is measured by the delivery of truly production-ready software.
VII-C Deployment: From Ethical Quagmires to Responsible Collaboration
The deployment of autonomous agents introduces a host of ethical, legal, and resource-related challenges that create an accountability vacuum and threaten sustainable adoption. Intellectual property frameworks are unprepared to resolve attribution for code co-created by humans, open-source libraries, and AI, creating significant legal uncertainty. Operationally, the exponential scaling of computational costs presents an unsustainable model, while the workforce faces potential deskilling. Most critically, when AI-generated code fails, the question of liability is diffused among the developer who prompted the agent, the organization that deployed it, and the creators of the model, creating a trilemma that obstructs adoption in mission-critical industries.
Navigating these challenges requires a concerted research effort toward responsible, sustainable, and collaborative AI systems. To establish accountability, agents must be built with inherent mechanisms for transparency, such as the ability to generate verifiable audit trails that explain design rationale and trace code to its sources. To ensure sustainability, research must prioritize resource-frugal agent architectures and cost-aware planning algorithms. Finally, the paradigm must shift from pure automation to human-centric augmentation, framing AI as a scaffolding tool that empowers and upskills developers. By designing agents as accountable, sustainable, and collaborative partners, we can build the foundation of trust required for their ethical and widespread adoption in the software engineering ecosystem.
VIII Challenges and Future Directions
@@
VIII-A Scalability: From Token Limits to Hierarchical Cognition
@@
VIII-B Evaluation: From Functional Correctness to Production Readiness
@@
VIII-C Domain Adaptation: From General Patterns to Specialized Expertise
@@
VIII-D Multi-Agent Coordination: From Individual Capability to Collective Intelligence
@@
VIII-E Continuous Learning: From Static Models to Evolving Expertise
@@
VIII-F Responsible Deployment: From Ethical Quagmires to Trustworthy Collaboration